OpenStudy (ohohaye):

How many zeros does the function have over the set of complex numbers? Function: x^4+x^3-6x^2-14x-12=0

Parth (parthkohli):

fundamental theorem of algebra

Parth (parthkohli):

do you know that theorem?

OpenStudy (denisaboichuk):

just one

OpenStudy (ohohaye):

Is it \[\frac{b \pm \sqrt{b ^{2}-4(a)(c)} }{ 2(a) }\]

OpenStudy (ohohaye):

@ParthKohli

Parth (parthkohli):

hehe no... this is not a quadratic equation the quadratic formula is a little wrong the fundamental theorem of algebra is not even remotely close

OpenStudy (ohohaye):

oh...

OpenStudy (ohohaye):

Oops...

Parth (parthkohli):

I think you can Google the theorem and it's just a second's work to find the number of complex solutions.

OpenStudy (ohohaye):

So it is where the graph passes the x-axis?

OpenStudy (ohohaye):

The roots are the zeros?

OpenStudy (ohohaye):

So are there 3 possible zeros/roots?

ganeshie8 (ganeshie8):

The peoblem is not about x intercepts. It is asking how many solutions are there to the given equation over complex numbers

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

Look at the given polynomial what is the degree ?

OpenStudy (ohohaye):

5

ganeshie8 (ganeshie8):

how did u get 5 ?

OpenStudy (ohohaye):

Do you mean on my problem?

ganeshie8 (ganeshie8):

Oh yes, your problem...

OpenStudy (ohohaye):

4

ganeshie8 (ganeshie8):

degree is 4, so there will be exactly 4 zeroes

OpenStudy (ohohaye):

Oh, so the leading degree is the amount of zeros?

ganeshie8 (ganeshie8):

Exactly! If the degree is 10, then there will be 10 zeroes

ganeshie8 (ganeshie8):

as simple as that

ganeshie8 (ganeshie8):

but you need to be careful, that count includes multiplicity of the zeroes too

OpenStudy (ohohaye):

That is when there are two numbers with the same degrees right?

ganeshie8 (ganeshie8):

Nope. consider below polynomial : \[p(x) = x^2-2x+1\]

ganeshie8 (ganeshie8):

In light of previous problem, can you tell the number of zeroes of above polynomial ?

OpenStudy (ohohaye):

Oh wait, I remember. That would be 3 because the 2x has an invisible 1 exponent right?

OpenStudy (astrophysics):

Note the degree!

OpenStudy (ohohaye):

its 2

ganeshie8 (ganeshie8):

right, so there will be exactly 2 zeroes

ganeshie8 (ganeshie8):

can you find them ?

OpenStudy (ohohaye):

x^2

OpenStudy (ohohaye):

Can you please just tell me

ganeshie8 (ganeshie8):

Incase if you just want the answer for your actual problem, the answer is 4

OpenStudy (ohohaye):

ok, so the number at the end has a coefficient of 0 and those are considered complex numbers?

ganeshie8 (ganeshie8):

you're asking questions with different levels of difficulty... by any chance are you skipping lessons ?

OpenStudy (astrophysics):

The link I sent earlier has all this info

OpenStudy (astrophysics):

Maybe just read it, has examples to not too long either https://www.mathsisfun.com/algebra/fundamental-theorem-algebra.html

ganeshie8 (ganeshie8):

its not wise to attempt homework with out reading the required material i can help you if you have a specific question... start somewhere :)

OpenStudy (ohohaye):

This is material that I went over at the beginning of the year and I forgot how to do it